OpenStudy (anonymous):

2 Algebra questions! I will award medal! 11. Which of the following expressions is true? (1 point) 4^3 • 4^4 = 4^12 5^2 • 5^3 > 5^5 3^2 • 3^5 = 3^10 5^2 • 5^4 < 5^8 12. Which of the following expressions is true? (1 point) 8^3 • 8^2 < 8^4 4^4 • 4^4 = 4^16 2^2 • 2^6 < 2^8 3^3 • 3^3 > 3^5

OpenStudy (anonymous):

When dealing with multiplying exponents, you essentially are adding the exponents together. Since, for example, 4^2 is just 4*4, thus (4^2)*(4^3) would be (4*4)*(4*4*4)

OpenStudy (anonymous):

I know how to figure out exponents... I just dont know how to add them. @stedan18

OpenStudy (anonymous):

Well, it's as simple as adding the numbers in their exponents. In my previous example, (4^2)*(4^3) = (4^5) | I am just adding the exponents (2 + 3 = 5) Now how could you apply that to your problem?

OpenStudy (anonymous):

11. Would be B. Because 5^2 + 5^3 = 5^5 ... right?

OpenStudy (anonymous):

@stedan18

OpenStudy (anonymous):

Try not to think of it as adding like that, just that you are adding the parts that are in the exponent. But, yes that is correct, but because (5^2)*(5^3) = (5*5)*(5*5*5) = (5*5*5*5*5) = (5^5)

OpenStudy (anonymous):

Ok! And 12 would be C, correct? @stedan18

OpenStudy (anonymous):

Pay close attention to the <, >, and = signs

OpenStudy (anonymous):

Oh! So its A! I didnt even notice the signs! :/ @stedan18

OpenStudy (anonymous):

Haha it's a very common mistake, I make it myself sometimes. However, you still don't quite have it. (8^3)*(8^2) would be 8^5, which is greater than 8^4

OpenStudy (anonymous):

Exactly... so A would be correct... thats what i said..

OpenStudy (anonymous):

@stedan18

OpenStudy (anonymous):

The < sign would indicate that A is stating that (8^5) is less than (8^4), which is incorrect.

OpenStudy (anonymous):

Oh, ma bad. the answer is D. @stedan18

OpenStudy (anonymous):

Exactly! Because you will have (3^3)*(3^3) >(3^5) which would lead to (3^6) > (3^5) which is true! Those lesser than and greater than signs will get you! haha